OpenStudy (anonymous):

Can you help me with: 3/5 < (3 - 2x)/-4 < 4/5

OpenStudy (anonymous):

Solve for x :

OpenStudy (anonymous):

@Hero @lgbasallote @.Sam. @Shane_B @eseidl @pfenn1 @sami-21 Please help me, thank you <3

OpenStudy (anonymous):

\[3/5 < \frac{ 3 - 2x }{ -4 } < 4/5\]

OpenStudy (pfenn1):

Try solving for two equations:\[\frac{3-2x}{-4}>\frac 35\]\[\frac{3-2x}{-4}<\frac45\]

OpenStudy (pfenn1):

I meant inequalities, not equations. Note:\[\frac{3-2x}{-4}=\frac{2x-3}{4}\]

OpenStudy (pfenn1):

Can you solve it from here?